# Measurement-only dynamical phase transition of topological and boundary orders in toric code and gauge-Higgs models

### v1

@@ IV Dynamics of boundary order parameter
Finally in Fig. 3(iii), we observe $p_{x}$-dependence of the late time value of $\chi_{SG_{z}}$ and $\chi_{SG_{x}}$ for $p_{s}=0.5$. For small $p_{x}$, the late time SG order is enhanced. Close look at behavior of the SG order signifies that a crossover, instead of a phase transition, takes place, i.e., the late time values of $\chi_{SGz}(\chi_{SGx})$ smoothly increase with decreasing (increasing) $p_{x}$. Further, see the inset of Fig. 3(iii), the variances of $\chi_{SGz}$ and $\chi_{SGx}$ show no significant behavior, mono-increase or decrease. This result is in contrast to the observation for the pure $Z$-measurement protocol corresponding to the deconfined(toric code)-Higgs transition studied in the previous paper Ref. KOI2024 . No quantities are indicating a bulk thermodynamic phase transition for the Higgs-confined phase. We expect that this reflects to the behavior of the boundary SG order indicating a novel bulk-boundary correspondence. We explained in the previous study KOI2024 that a finite expectation value of open Wilson loops (EVOWL) generates effective spin couplings on the rough boundary and the SG order emerges as a result. On the deconfined(toric code)-Higgs phase transition, the EVWL (more precisely, Fredenhagen-Marcu string order) exhibits a transition-like behavior similar to the magnetization in the (2+1)D Ising model, whereas in the confined-Higgs crossover regime, both the Wilson and ’t Hooft string symmetries do not emerge as Fradkin-Shenker observation dictates (see also Xu2024 ). Reflecting this smooth change in the bulk properties, the SG order exhibits crossover-like behavior instead of a genuine transition as we observed numerically.
@@ V Dynamics of topological entanglement entropy
We move on to the dynamics of the TEE. We first study the recovery effect by the measurement of layer 2 for a fixed probability of the measurement of layer 1. The numerical results for various sizes $L_{x}$ are displayed in Fig. 5 for the most competitive case of layer 1, $p_{x}=p_{z}=0.5$. For a large recovery case $p_{s}=0.80$, the TEE starts with $\gamma=-1$ and exhibits almost no changes in the long time evolution. The stationary state sustains TO in the bulk staying in the deconfined phase. As shown in Fig. 5 (d), in the whole time evolution, fluctuations are rather small. On the other hand for a relatively weak recovery case with $p_{s}=0.50$, behavior of the TEE drastically changes from that of large recovery case as seen in Figs. 5 (a) and (b). The late time value of $\gamma$ saturates to $\gamma=0$, indicating that the stationary state loses the initial TO in the bulk. We also find that the fluctuation of the TEE is large in the regime where the TEE changes drastically. This behavior of the TEE in this MoC obviously indicates dynamic phase transition from the deconfnied to Higgs phases. In order to obtain a critical probability, $p^{c}_{s}$, for the transition, we further observe the late time value of the TEE around $t/L^{x}\sim 20$ and find $p^{c}_{s}\simeq 0.65$. The results are displayed in Figs. 5 (a) and (c). There, the saturation value of $\gamma$ is fluctuating around $\gamma=-0.5$ and its fluctuation exhibits large values for a long period, exhibiting a small system-size dependence.
@@
Figure 5: (a) Time evolution of topological entanglement entropy $\gamma$ for $p_{x}=p_{z}=0.5$ and various $p_{s}$ and system sizes. (b) $\sim$ (d): Fluctuation of topological entanglement entropy. We used $1000$, $1000$, $500$ and $500$ samples for $L_{x}=14,18,22,$ and $26$.
Finally, we summarize the saturation vale of $\gamma$ as a function of $p_{s}$ for various system sizes in Fig. 6. We find a clear step-function-like behavior of the TEE and a peak of the fluctuations of the TEE is located at $p_{s}\simeq 0.65$. We expect that in the thermodynamic limit, the saturation value of the TEE becomes a genuine step function of $p_{s}$. These results imply that there exists a clear bulk phase transition emerging by varying the recovery probability $p_{s}$, at which the bulk topological order vanishes.
We also investigate the behavior of the TEE for a non-contractible partition pattern of the system as in Ref. Levin_Wen , and obtain similar results to the above, which are shown in Appendix B. The initial TC state has $\gamma=-2$ for the non-contractible partition, and the state changes its value depending on the probability rates. The critical probability is estimated as $0.65$ for $p_{x}=p_{z}=0.5$.
Figure 6: Long time saturation values of topological entanglement entropy for $p_{x}=p_{z}=0.5$. Topological entanglement entropy and its fluctuation indicate that the bulk phase transition takes place at $p_{s}\simeq 0.65$. Numerical calculations are performed up to $t=20L$. We took both sample and time average. As for the time average, we employ the last nine data points. We used $1000$, $1000$, $500$, and $500$ samples for $L_{x}=14,18,22$ and $26$.
@@ VI Conclusion and discussion
Figure 7: Time evolution of the spin-glass (SG) order and topological entanglement entropy for $p_{x}=0.2$. Data of both the cases show that the topological entanglement entropy tends to vanish before the development of the SG order. This behavior can be interpreted that condensation of Wilson loop in the bulk is the origin of the SG order on the boundary. The sample averages are taken over $200$ samples for both $L_{x}=14$ and $L_{x}=18$.
The present work clarified the dynamics of MoC, in which we performed competitive measurements on the TC state. There, the measurement operators correspond to the terms in the Hamiltonian of the TC model in magnetic fields. In this MoC, stationary states, which seem to have properties of Higgs, confined and critical phases, emerge by the ‘time evolution’ from the exact stabilizer state of the deconfined (toric code) phase on the cylinder geometry. We identified physical properties of emergent states by numerically observing the boundary long-range magnetic orders and TEE in the bulk. We found that the critical measurement probability ratios corresponds to the ratios of the parameters in the lattice gauge-Higgs model Hamiltonian.
In the present MoC, we observed simultaneously both the SG LRO on the boundary and TEE in the bulk. This gives us an important insight into the relation between the SPT and intrinsic TO in the gauge-Higgs models. As shown in Fig. 7, the ‘time evolution’ of the SG order on the boundaries and TEE in the bulk is observed clearly to find that the bulk topological transition takes place first and the SG transition follows that. This behavior is observed generally. As we explained in the main text, the SG order stems from the bulk one-form-symmetry, in particular, condensation of the Wilson (’t Hooft) string. This condensate induces effective Ising-type long-range couplings between spins on the boundaries producing the SG order. Numerical study in Fig. 7 can be interpreted naturally in such a way that a sufficiently large coupling between boundary spins is necessary for the SG order to emerge. In fact, the recent study on Wilson string condensate for lattice gauge-Higgs model in Ref. Xu2023 shows that the condensate behaves as an order parameter and exhibits $(2+1)-D$ Ising spin criticality developing continuously from zero. The difference in the location of the two transitions, which was observed already in our previous work KOI2024 , supports the above observation concerning to the bulk-boundary correspondence.
@@ Acknowledgements
Appendix A: Non-local gauge invariant operator
@@
To qualitatively understand the boundary LRO in the Higgs phase in the toric code system with the upper rough and lower smooth boundaries on the cylinder geometry, we introduce non-local gauge invariant operator (NLGIO) symmetries for $H_{\rm TC}$ with $J_{x}=0$. Here, we focus on the upper rough boundary and Higgs phase with noticing that parallel argument is possible for the lower smooth boundary and confined phase. Then, the symmetry has two types: (I) $G_{lo,1}\equiv\prod_{\ell\in\Gamma_{0}}\sigma^{z}_{\ell}$ with an arbitrary close loop $\Gamma_{0}$, (II) $G_{lo,2}=\sigma^{z}_{\ell_{r1}}\biggl{[}\prod_{\ell\in\Gamma_{b}}\sigma^{z}_{\ell}\biggr{]}\sigma^{z}_{\ell_{r2}}$ where $\ell_{r1}$ and $\ell_{r2}$ are two arbitrary dangling links on the upper rough boundary and a string in the bulk $\Gamma_{b}$ connecting the $\ell_{r1}$ and $\ell_{r2}$ links. Both these operators satisfy $[H_{\rm TC},G_{lo,1(2)}]=0$ for $J_{x}=0$. The second-type NLGIO $G_{lo,2}$ plays a key role to understand the boundary LRO through the Higgs condensation. That is, the Higgs phase can be regarded as a symmetry-restored state of the second-type NLGIO $G_{lo,2}$. In the deep Higgs phase, the local $\sigma^{z}_{\ell}$ stabilizer generator is proliferated in the bulk, leading to a finite string order, $\langle\prod_{\Gamma_{b}}\sigma^{z}_{\ell}\rangle\neq 0$ in the system, regarded as the Higgs condensation. As the state of the Higgs phase $|\psi_{\rm Higgs}\rangle$ respects the NLGIO symmetry, then the following relation is satisfied, $G_{lo,2}|\psi_{\rm Higgs}\rangle=\sigma^{z}_{\ell_{r1}}\biggl{[}\prod_{\ell\in\Gamma_{b}}\sigma^{z}_{\ell}\biggr{]}\sigma^{z}_{\ell_{r2}}|\psi_{\rm Higgs}\rangle\propto|\psi_{\rm Higgs}\rangle$. That is, the Higgs condensation in the bulk gives the result $\langle\sigma^{z}_{\ell_{r1}}\sigma^{z}_{\ell_{r2}}\rangle\neq 0$, the emergence of the LRO on the rough boundary. We expect that the NLGIO symmetries have a similar robustness to the general one-form-symmetry, and the above observation holds for a finite-$J_{x}$ system as long as it belongs to the Higgs phase. Numerical studies in the present work support this expectation.
@@
Appendix B: topological entanglement entropy of another partition
Figure 9: Time evolution of topological entanglement entropy Levin_Wen for $p_{x}=p_{z}=0.5$ and various $ps$ and system sizes $L_{x}$. We fixed $L_{y}=10$. Panels (b)-(d) represent fluctuations of TEE for (b) $p_{s}=0.50$, (b) $p_{s}=0.65$, and (c) $p_{s}=0.80$, respectively. The sample averages are taken over $1000,1000,$ and $500$ samples for both $L_{x}=14,18$ and $22$.

### v2

@@ IV Dynamics of boundary order parameter
Finally in Fig. 3(iii), we observe $p_{x}$-dependence of the late time value of $\chi_{SG_{z}}$ and $\chi_{SG_{x}}$ for $p_{s}=0.5$. For small $p_{x}$, the late time SG order is enhanced. Close look at behavior of the SG order signifies that a crossover, instead of a phase transition, takes place, i.e., the late time values of $\chi_{SGz}(\chi_{SGx})$ smoothly increase with decreasing (increasing) $p_{x}$. To ensure if the phase transition exists, we further focus on the variances of $\chi_{SGz}$ and $\chi_{SGx}$. The inset of Fig. 3(iii) shows the variances of $\chi_{SGz}$ and $\chi_{SGx}$, and we found that these two quantities do not show any significant behavior, i.e., monotonic increase or decrease. This result is in contrast to the observation for the pure $\sigma^{z}$-measurement protocol corresponding to the deconfined (toric code)-Higgs transition studied in the previous paper Ref. KOI2024 , where $\sigma(\chi_{SG_{z}})/L_{x}$ shows a sharp peak, and its peak becomes more significant as $L_{x}$ increases, implying the existence of the conventional second-order phase transition. In the present case, however, no quantities are indicating a bulk thermodynamic phase transition for the Higgs-confined phase. We expect that this reflects to the behavior of the boundary SG order indicating a novel bulk-boundary correspondence. We explained in the previous study KOI2024 that a finite expectation value of open Wilson loops (EVOWL) generates effective spin couplings on the rough boundary and the SG order emerges as a result. On the deconfined (toric code)-Higgs phase transition, the EVWL (more precisely, Fredenhagen-Marcu string order) exhibits a transition-like behavior similar to the magnetization in the (2+1)D Ising model, whereas in the confined-Higgs crossover regime, both the Wilson and ’t Hooft string symmetries do not emerge as Fradkin-Shenker observation dictates (see also Xu2024 ). Reflecting this smooth change in the bulk properties, the SG order exhibits crossover-like behavior instead of a genuine transition as we observed numerically. This observation obtained by the numerical study on the boundary SG order is complementary to the work in Ref. Verresen2022 (in which the boundary transition between Higgs-confined regimes was investigated by density-matrix-renormalization-group methods).
@@ V Dynamics of topological entanglement entropy
We move on to the dynamics of the TEE. We first study the recovery effect by the measurement of layer 2 for a fixed probability of the measurement of layer 1. The numerical results for various sizes $L_{x}$ are displayed in Fig. 5 for the most competitive case of layer 1, $p_{x}=p_{z}=0.5$. For a large recovery case $p_{s}=0.80$, the TEE starts with $\gamma=-1$ and exhibits almost no changes in the long time evolution. The stationary state sustains TO in the bulk staying in the deconfined phase. As shown in Fig. 5 (d), in the whole time evolution, fluctuations are rather small. On the other hand for a relatively weak recovery case with $p_{s}=0.50$, behavior of the TEE drastically changes from that of large recovery case as seen in Figs. 5 (a) and (b). The late time value of $\gamma$ saturates to $\gamma=0$, indicating that the stationary state loses the initial TO in the bulk. We also find that the fluctuation of the TEE is large in the regime where the TEE changes drastically. This behavior of the TEE in this MoC obviously indicates dynamic phase transition from the deconfnied to Higgs phases. In order to obtain a critical probability, $p^{c}_{s}$, for the transition, we further observe the late time value of the TEE around $t/L=20$ (for $p_{s}=0.5,0.6,0.75,0.8,0.9$) and $t/L=30$ (for $p_{s}=0.65,0.7$) and find $p^{c}_{s}\simeq 0.65$. The results are displayed in Figs. 5 (a) and (c). There, the saturation value of $\gamma$ is fluctuating around $\gamma=-0.5$ and its fluctuation exhibits large values for a long period, exhibiting a small system-size dependence.
@@
Figure 5: (a) Time evolution of topological entanglement entropy $\gamma$ for $p_{x}=p_{z}=0.5$ and various $p_{s}$ and system sizes. (b) $\sim$ (d): Fluctuation of topological entanglement entropy. We used $2000$, $2000$, $1000$ and $1000$ samples for $L_{x}=14,18,22$ and $26$, respectively.
Finally, we summarize the saturation vale of $\gamma$ as a function of $p_{s}$ for various system sizes in Fig. 6. We find a clear step-function-like behavior of the TEE (panel (a)) and a peak of the fluctuations of the TEE is located at $p_{s}\simeq 0.65$ (panel (b)). We expect that in the thermodynamic limit, the saturation value of the TEE becomes a genuine step function of $p_{s}$. These results imply that there exists a clear bulk phase transition emerging by varying the recovery probability $p_{s}$, at which the bulk topological order vanishes.
We also investigate the behavior of the TEE for a non-contractible partition pattern of the system as in Ref. Levin_Wen , and obtain similar results to the above, which are shown in Appendix C. The initial TC state has $\gamma=-2$ for the non-contractible partition, and the state changes its value depending on the probability rates. The critical probability is estimated as $0.65$ for $p_{x}=p_{z}=0.5$.
Figure 6: Long time saturation values of topological entanglement entropy for $p_{x}=p_{z}=0.5$. Topological entanglement entropy and its fluctuation indicate that the bulk phase transition takes place at $p_{s}\simeq 0.65$. Numerical calculations are performed up to $t=20L$ and $t=30L$ for $p_{s}=0.5,0.6,0.75,0.8,0.9$ and $p_{s}=0.65,0.70$, respectively. We took both sample and time average. As for the time average, we employ the last nine data points. We used $2000$, $2000$, $1000$, and $1000$ samples for $L_{x}=14,18,22$ and $26$, respectively.
@@ VI Conclusion and discussion
The present work clarified the dynamics of MoC, in which we performed competitive measurements on the TC state on a cylinder geometry with distinct upper and lower boundary conditions to observe the duality property between the Higgs and confined phases. There, the measurement operators correspond to the terms in the Hamiltonian of the TC model in magnetic fields. In this MoC, stationary states, which seem to have properties of Higgs, confined and critical phases, emerge by the ‘time evolution’ from the exact stabilizer state of the deconfined (toric code) phase on the cylinder geometry. We identified physical properties of emergent states by numerically observing the boundary long-range magnetic orders and TEE in the bulk. We found that the critical measurement probability ratios corresponds to the ratios of the parameters in the lattice gauge-Higgs model Hamiltonian. We further investigated how the initial TC state develops under various settings of the measurement. In particular, we observed that the Higgs and confined phases are separated from the deconfined phase by topological entanglement entropy, whereas the Higgs and confined phases are distinguishable with each other via a crossover of boundary modes.
Figure 7: Time evolution of the spin-glass (SG) order and topological entanglement entropy for $p_{x}=0.2$. Data of both the cases show that the topological entanglement entropy tends to vanish before the development of the SG order. This behavior can be interpreted that condensation of Wilson loop in the bulk is the origin of the SG order on the boundary. The sample averages are taken over $1000$ samples for both $L_{x}=14$ and $L_{x}=18$.
In the present MoC, we observed simultaneously both the SG LROs on the upper-rough and lower-smooth boundaries and TEE in the bulk. This gives us an important insight into the relation between the SPT and intrinsic TO in the gauge-Higgs models. As shown in Fig. 7, the ‘time evolution’ of the SG order on the boundaries and TEE in the bulk is observed clearly to find that the bulk topological transition takes place first and the SG transition follows that. This behavior is observed generally. As we explained in the main text, the SG order stems from the bulk one-form-symmetry, in particular, condensation of the Wilson (’t Hooft) string. This condensate induces effective Ising-type long-range couplings between spins on the boundaries producing the SG order. Numerical study in Fig. 7 can be interpreted naturally in such a way that a sufficiently large coupling between boundary spins is necessary for the SG order to emerge. In fact, the recent study on Wilson string condensate for lattice gauge-Higgs model in Ref. Xu2024 shows that the condensate behaves as an order parameter and exhibits $(2+1)-D$ Ising spin criticality developing continuously from zero. The difference in the location of the two transitions, which was observed already in our previous work KOI2024 , supports the above observation concerning to the bulk-boundary correspondence.
@@ Acknowledgements
Appendix A: Parameter ratio and probability ratio correspondence
In the previous study by two of the authors of this work KI2023 , a conjecture was given; parameter ratio and probability ratio correspondence. The target Hamiltonian is the following,
$H_{\rm stab}=-\sum^{L-1}_{j=0}\sum^{M}_{\alpha=1}J^{\alpha}K^{\alpha}_{j},$
where $J^{\alpha}(>0)$ and $L$ is the total number of sites $\{j\}$, $\alpha$ represents $M$-types of stabilizers anti-commuting with each other, i.e., $\{K^{\alpha}_{j}\}$ satisfy $[K^{\alpha}_{j},K^{\alpha}_{k}]=0$ and $(K^{\alpha}_{j})^{2}=1$, and for different types of stabilizers, $[K^{\alpha}_{j},K^{\beta}_{k}]\neq 0$ and $\{K^{\alpha}_{j},K^{\beta}_{k}\}=0$ ($\alpha\neq\beta$).
To construct a corresponding MoC, we pick up a single stabilizer among the different types of $K^{\alpha}_{j_{0}}$ with a probability $p^{\alpha}$ and choose a target site $j_{0}$ with equal probability $1/L$ at each time step. We set the probability condition of the choice of the type of the stabilizer, such as $\sum_{\alpha}p^{\alpha}=1$. The setup is the same as that of the previous works Lavasani2021 ; Klocke2022 .
Then, the ground-state phase diagram of $H_{\rm stab}$ determined by the parameters $J^{\alpha}$ is very close to the phase diagram of the steady state in the MoC, which is determined by an ensemble average of the measurement pattern of the MoC. This fact indicates that the ratio of parameters $\frac{J^{\alpha}}{J^{\beta}}$ is related to the ratio of probabilities $\frac{p^{\alpha}}{p^{\beta}}$, that is, $\frac{J^{\alpha}}{J^{\beta}}\longleftrightarrow\frac{p^{\alpha}}{p^{\beta}}$. This relationship is nothing but the explicit form of “parameter ratio-probability ratio correspondence”. The qualitative verification of the above consideration was analytically carried out in Ref. KI2023 .
Appendix B: Non-local gauge invariant operator
@@
To qualitatively understand the boundary LRO in the Higgs phase in the toric code system with the upper rough and lower smooth boundaries on the cylinder geometry, we introduce non-local gauge invariant operator (NLGIO) symmetries for $H_{\rm TC}$ with $J_{x}=0$. Here, we focus on the upper rough boundary and Higgs phase with noticing that parallel argument is possible for the lower smooth boundary and confined phase. Then, the symmetry has two types: (I) $G_{lo,1}\equiv\prod_{\ell\in\Gamma_{0}}\sigma^{z}_{\ell}$ with an arbitrary close loop $\Gamma_{0}$, (II) $G_{lo,2}=\sigma^{z}_{\ell_{r1}}\biggl{[}\prod_{\ell\in\Gamma_{b}}\sigma^{z}_{\ell}\biggr{]}\sigma^{z}_{\ell_{r2}}$ where $\ell_{r1}$ and $\ell_{r2}$ are two arbitrary dangling links on the upper rough boundary and a string in the bulk $\Gamma_{b}$ connecting the $\ell_{r1}$ and $\ell_{r2}$ links.
Both these operators satisfy $[H_{\rm TC},G_{lo,1(2)}]=0$ for $J_{x}=0$. The second-type NLGIO $G_{lo,2}$ plays a key role to understand the boundary LRO through the Higgs condensation. That is, the Higgs phase can be regarded as a symmetry-restored state of the second-type NLGIO $G_{lo,2}$.
Figure 9: Time evolution of topological entanglement entropy Levin_Wen for $p_{x}=p_{z}=0.5$ and various $ps$ and system sizes $L_{x}$. We fixed $L_{y}=10$. Panels (b)-(d) represent fluctuations of TEE for (b) $p_{s}=0.50$, (b) $p_{s}=0.65$, and (c) $p_{s}=0.80$, respectively. The sample averages are taken over $1000$, $1000$, and $500$ samples for both $L_{x}=14,18$ and $22$, respectively.
In the deep Higgs phase, the local $\sigma^{z}_{\ell}$ stabilizer generator is proliferated in the bulk, leading to a finite string order, $\langle\prod_{\Gamma_{b}}\sigma^{z}_{\ell}\rangle\neq 0$ in the system, regarded as the Higgs condensation. As the state of the Higgs phase $|\psi_{\rm Higgs}\rangle$ respects the NLGIO symmetry, then the following relation is satisfied, $G_{lo,2}|\psi_{\rm Higgs}\rangle=\sigma^{z}_{\ell_{r1}}\biggl{[}\prod_{\ell\in\Gamma_{b}}\sigma^{z}_{\ell}\biggr{]}\sigma^{z}_{\ell_{r2}}|\psi_{\rm Higgs}\rangle\propto|\psi_{\rm Higgs}\rangle$. That is, the Higgs condensation in the bulk gives the result $\langle\sigma^{z}_{\ell_{r1}}\sigma^{z}_{\ell_{r2}}\rangle\neq 0$, the emergence of the LRO on the rough boundary. We expect that the NLGIO symmetries have a similar robustness to the general one-form-symmetry, and the above observation holds for a finite-$J_{x}$ system as long as it belongs to the Higgs phase. Numerical studies in the present work support this expectation.
@@
Appendix C: topological entanglement entropy of another partition